A. It takes 1.0 kg of coal to produce 28.8x10^6 Joules of heat. If a power plant uses 16.5 kg of coal per second, what is the electrical power output in Watts or Megawatts?

B. What mass of water uses 28.8x10^6 Joules of heat to go from 308K to 853K?
Answer #1

A ) Use equation,

Power = Total Energy/ time =  (no of kg used *energy per kg)/time = no of kg used *(energy per kg/time) = 16.5kg*(28.8*10^6 J/kg)/s = 4.75*10^8 watt

B) Use equation,

Q= mc(Tf-Ti)

28.8*10^6  = m*4186*(853-308)     => m = 12.62 kg